# From CLIP to DINO: Visual Encoders Shout in Multi-modal Large Language Models

### v1

@@ 5 Experiments
Training Details. Similar to previous MLLM methods, COMM is trained in two stages. In the first pretraining stage, we train the model on the reorganized vision-language dataset as (Chen et al., 2023), including public VQA, Image Captioning datset and several datasets containing positional annotation RefCOCO, visual gemone (Krishna et al., 2017) and Visual-7W (Mani et al., 2020). The first pretraining stage is conducted for 100K steps. In the second instruction tuning stage, we set the sampling ratio to 50% on LLaVA-Instruct-150K (Liu et al., 2023b) and Shikra-RD (Chen et al., 2023). Instead of 224 × 224 resolution currently used by existing MLLMs, we use 336 × 336 resolution to reduce the information loss caused by image down-sampling and promote the fine-grained perception ability. In both stages, we freeze the visual encoder and tune all parameters in LLMs, alignment layer and multi-level feature fusion module. We adopt AdamW (Loshchilov & Hutter, 2019) as the optimizer and cosine annealing scheduler (Loshchilov & Hutter, 2017) as learning rate scheduler with an initial learning rate of 2e-5 and global batch size of 64. All training runs on 8 NVIDIA A800 GPUs. It takes around 100h for stage one training and 20h for stage two.
Table 2: Results on standard referring expression comprehension (REC) task. Generalist VL models can perform various vision-language tasks. Specialist models are designed specifically for localization tasks or generalist pretraining models that have undergone finetuning. The results of Shikra and Qwen are taken from their papers.
@@ 5.1 Referring Expression Comprehension
To evaluate the fine-grained understanding and positioning capability of our model, we investigate the referring expression comprehension task on benchmarks as RefCOCO (Kazemzadeh et al., 2014), RefCOCO+ (Mao et al., 2016) and RefCOCOg (Mao et al., 2016), where models are asked to localize the object described with an expression. As shown in Table 2, compared with generalist VL models and previous SOTA MLLMs, COMM achieves significant performance gain on all benchmarks, i.e., COMM-7B outperforms Shikra-13B and Qwen-VL-7B-Chat by 4.87% and 3.10% accuracy on average, respectively. With more powerful visual capabilities of our proposed fusion model, we can evidently surpass recent SOTA MLLMs in a more efficient way, e.g., using a smaller LLM than Shikra (7B vs. 13B) and less training data than Qwen (3.6M vs. 1.4B). Besides, our generalist model even achieves comparable results with specialist SOTA methods, showing the superior grounding ability of our MLLMs.
@@ 5.2 Referring Expression Generation
Table 4: Object hallucination benchmark using POPE evaluation pipeline (Li et al., 2023d). The results of Shikra-7B are taken from its paper. Except for Shikra-7B, the other results are obtained from (Li et al., 2023d).
@@
Table 5: Results on visual question answering (VQA) and image captioning. For VQA, we evaluate SOTA generalist models and our COMM onVQAv2 and OK-VQA following the normalization rules. Shikra and LLaVA-1.5 (Liu et al., 2023a) is based on the 13B variant. For image captioning, we evaluate them on COCO and Flickr30k in CIDEr score. We call Flamingo as FM for short.
@@ 5.3 Object Hallucination Benchmark
We compare our model against the baseline models on the hallucination evaluation dataset recently introduced by POPE (Li et al., 2023d), which randomly selects 500 images from COCO (Caesar et al., 2018). Table 4 shows that COMM surpasses recent popular MLLMs with 1.44% and 4.95% higher accuracy on average than Shikra and InstrutBLIP, respectively. By enhancing the fine-grained visual capabilities, COMM can effectively alleviate the object hallucination problem.
@@ 5.4 Visual Question Answering and Image Captioning
We evaluate COMM on conventional VL tasks of VQA and Image Captioning. Specifically, image captioning requires the model to generate description for the given image and VQA asks the model to generate answer for the given image-question pair. For image captioning, we choose COCO (Chen et al., 2015) and Flickr30K (Plummer et al., 2015) as benchmarks and report the CIDEr score. For VQA task, we experiment on VQAv2 (Antol et al., 2015) and OK-VQA (Marino et al., 2019). As shown in Table 5, COMM achieves state-of-the-art performance on image captioning task, i.e., 88.2 CIDEr score on Flickr30K and 132.7 CIDEr score on COCO, even outperforms previous SOTA models with much more parameters (e.g., Shikra-13B with 13B parameters) or much more training data (e.g., Qwen with 1.4B data). For VQA task, our model also shows significant advantages compared to other MLLMs. On VQAv2 val, dev and std, our model achieves 79.05, 81.04 and 81.17 accuracy respectively, which surpasses recent proposed Shikra with the same training data and procedure by a large margin, demonstrating the effectiveness of merging visual embeddings of DINOv2 and CLIP for enhancing visual capabilities. Besides, our COMM model outperforms Qwen with 1.54 and 0.58 accuracy gain on VQAv2 dev and OK-VQA respectively with less VQA training data, i.e., we use 0.6M and Qwen with 3.6M. Training with more VQA data might further improve performance and we leave it as future work.
5.5 Demonstrations
As shown in Fig. 5, our COMM model exhibits a multitude of promising capabilities including visual grounding, fine-grained region understanding and robustness to object hallucination. The first example showcases our strong fine-grained perception ability, which identifies implicit strawberries in a blender. The second example exhibits our strong visual grounding ability to successfully locates the object of sugar. The third case demonstrates our robustness to object hallucination. In contrast, Shikra fails on these challenging cases, showing the superior capabilities of our model.
6 Conclusion
Acknowledgements
We thank Weihang Li, Xinghao Wang and Junjie Wang for valuable discussions on the experimental evaluation.
Appendix
Appendix A Analysis of DeiT as the Visual Branch
We conduct experiment with recent supervised learning model DeiT III (Touvron et al., 2022). Since the visual model does not align with text, we employ a non-linear MLP layer as the alignment network. Here we use the ViT-Large variant as the visual branch and employ Vicuna-7B as the language branch of MLLMs. The architecture and training process follow Shikra (Chen et al., 2023) but with fewer training iterations (9400 iterations, batch size 16 on 4 A800) to save the computation cost. As shown in Table 6, DeiT suffers from evident performance degradation as the visual branch of MLLMs. We speculate that maybe the supervised training is so strong that it learns specialized visual space, making it difficult to align with the word embedding space.
Table 6: Comparison of the visual model using CLIP, DINOv2 with our proposed multi-level feature merging (MFM), MAE and DeiT. Experiments are conducted on referring expression comprehension and object hallucination benchmark on Random (R), Adversarial (A), and Popular (P). MAE-20 denotes using the features output by the 20-th layer of MAE. DeiT-20 denotes using the features output by the 20-th layer of DeiT.
Visual Model RefCOCO+ RefCOCOg RefCOCO POPE
test-A test-B val test-u val-u test-A test-B val A/P/R
CLIP w/ MFM 73.7 53.8 64.3 69.1 70.3 83.8 68.4 76.4 80.7/84.2/85.8
DINOv2 w/ MFM 75.3 59.3 67.0 73.0 71.8 84.4 74.1 79.6 80.3/84.2/85.5
MAE-20 64.7 49.4 56.8 63.7 62.8 77.9 68.6 73.6 66.8/71.1/76.7
MAE-22 65.9 50.0 58.5 64.2 63.2 79.3 69.8 74.9 68.0/71.2/77.5
MAE-24 67.7 51.5 60.2 65.4 64.3 80.5 69.9 75.8 69.8/76.1/80.8
DeiT-20 18.4 13.0 15.9 17.0 16.2 29.0 21.6 25.7 66.2/69.6/77.9
DeiT-22 25.3 15.4 19.4 22.6 21.8 36.9 25.3 32.0 67.9/71.6/78.7
Appendix B Ablation on the MLP of DINOv2
@@
We conduct ablation study on the MLP module of DINOv2 for aligning visual and text embedding space. Here we use the ViT-Large variant as the visual branch and employ Vicuna-7B as the language branch of MLLMs. The architecture and training process follow Shikra (Chen et al., 2023) but with fewer training iterations (9400 iterations, batch size 16 on 4 A800) to save the computation cost. We ablate on the number and the expanding ratio of MLP module. Table 7 shows that increasing the number of MLP to 2 can evidently improve performance, demonstrate the effectiveness of using a more powerful network to align the vision only model DINOv2 to the word embedding space. However, increasing the number beyond 2 suffers the degraded performance. For the expanding ratio, increasing to 8 can improve performance, while increasing to 16 does not achieve significant performance gain. Moreover, we experiment with one linear layer, which suffers severe performance degradation. Thus, non-linear MLP is necessary for aligning the features of vision-only DINOv2 to the word embedding space.
@@
Appendix C Evaluation Benchmark for Multimodal Large Language Models
We evaluate our final model on the recently proposed MME benchmark to further demonstrate its strong generalizability to follow a diverse range of instructions. MME benchmark measures both perception and cognition abilities on a total of 14 subtasks. We report the average results of perception tasks and cognition tasks, respectively. As shown in Table 8, the performance of our COMM model outperforms Shikra by a large margin of 176.59 and 39.95 score for perception and cognition abilities on MME benchmark, demonstrating the effectiveness of enhancing visual capabilities with our fused visual embeddings. Compared with Qwen, our model suffers a litter performance degradation, and we speculate it is due to less training data than Qwen (3.6M vs. 1.4B) and we will try to train with more data for evaluation in the future work. Besides, our model has slightly lower perception score than LLaVA-1.5, since the latter model employs more powerful recent Vicuna v1.5 as the language model. We will build our model on more powerful language models in future work.
Table 8: Zero-shot evaluation of perception and cognition abilities on MME benchmark.
COMM-13B COMM-7B ${}^{\dagger}$ Qwen InstructBLIP LLaVA-1.5 MiniGPT-4 mPLUG-Owl
Perception 1421.00 1387.00 1210.41 1487.58 1212.82 1531.32 866.57 967.34
Cognition 315.00 314.60 274.65 360.71 291.79 295.36 292.14 276.07
Appendix D More Demonstrations
We provide additional demonstrations of our COMM model in this section to demonstrate a multitude of promising capabilities including visual grounding, fine-grained region understanding and robustness to object hallucination. For instance, we showcase Referring Expression Comprehension in Fig. 6 and Object Hallucination in Fig. 7.
@@ Appendix D More Demonstrations
We also include cases of Spot Captioning in Fig. 8 and Referring Expression Generation in Fig. 9.
Figure 8: Spotting Captioning using our COMM-7B. The task requires the model to describe the image and spots the mentioned objects or regions using points or boxes.
Figure 9: Referring Expression Generation (REG) using our COMM-7B. This task aims to generate a unique description for a specified location.

### v3

@@ 5 Experiments
Training Details. Similar to previous MLLM methods, COMM is trained in two stages. In the first pretraining stage, we train the model on the reorganized vision-language dataset as [8], including public VQA, Image Captioning datset and several datasets containing positional annotation RefCOCO, visual gemone [17] and Visual-7W [28]. The first pretraining stage is conducted for 100K steps. In the second instruction tuning stage, we set the sampling ratio to 50% on LLaVA-Instruct-150K [23] and Shikra-RD [8]. Instead of 224 × 224 resolution currently used by existing MLLMs, we use 336 × 336 resolution to reduce the information loss caused by image down-sampling and promote the fine-grained perception ability. In both stages, we freeze the visual encoder and tune all parameters in LLMs, alignment layer and multi-level feature fusion module. We adopt AdamW [27] as the optimizer and cosine annealing scheduler [26] as learning rate scheduler with an initial learning rate of 2e-5 and global batch size of 64. All training runs on 8 NVIDIA A800 GPUs. It takes around 100h for stage one training and 20h for stage two.
Table 2: Results on standard referring expression comprehension (REC) task. Generalist VL models can perform various vision-language tasks. Specialist models are designed specifically for localization tasks or generalist pretraining models that undergone finetuning. The results of Shikra, Qwen, Ferret and Griffon are from their papers.
@@ 5.1 Referring Expression Comprehension
To evaluate the fine-grained understanding and positioning capability of our model, we investigate the referring expression comprehension task on benchmarks as RefCOCO [15], RefCOCO+ [29] and RefCOCOg [29], where models are asked to localize the object described with an expression. As shown in Table 2, compared with generalist VL models and previous SOTA MLLMs, COMM achieves significant performance gain on all benchmarks, i.e., COMM-7B outperforms Shikra-13B and Qwen-VL-7B-Chat by 4.87% and 3.10% accuracy on average, respectively. With more powerful visual capabilities of our proposed fusion model, we can evidently surpass recent SOTA MLLMs in a more efficient way, e.g., using a smaller LLM than Shikra (7B vs. 13B) and less training data than Qwen (3.6M vs. 1.4B). Besides, our generalist model even achieves comparable results with specialist SOTA methods, showing the superior grounding ability of our MLLMs.
@@ 5.2 Referring Expression Generation
Table 4: Object hallucination benchmark using POPE evaluation pipeline [21]. The results of Shikra-7B are taken from its paper. Except for Shikra-7B, the other results are obtained from [21].
@@
Table 5: Results on visual question answering (VQA) and image captioning. For VQA, we evaluate SOTA generalist models and our COMM onVQAv2 and OK-VQA following the normalization rules. Shikra and LLaVA-1.5 [22] is based on the 13B variant. For image captioning, we evaluate them on COCO and Flickr30k in CIDEr score. We call Flamingo as FM for short.
@@ 5.3 Object Hallucination Benchmark
We compare our model against the baseline models on the hallucination evaluation dataset recently introduced by POPE [21], which randomly selects 500 images from COCO [6]. Table 4 shows that COMM surpasses recent popular MLLMs with 1.44% and 4.95% higher accuracy on average than Shikra and InstrutBLIP, respectively. By enhancing the fine-grained visual capabilities, COMM can effectively alleviate the object hallucination problem.
@@ 5.4 Visual Question Answering and Image Captioning
We evaluate COMM on conventional VL tasks of VQA and Image Captioning. Specifically, image captioning requires the model to generate description for the given image and VQA asks the model to generate answer for the given image-question pair. For image captioning, we choose COCO [9] and Flickr30K [35] as benchmarks and report the CIDEr score. For VQA task, we experiment on VQAv2 [2] and OK-VQA [30]. As shown in Table 5, COMM achieves state-of-the-art performance on image captioning task, i.e., 88.2 CIDEr score on Flickr30K and 132.7 CIDEr score on COCO, even outperforms previous SOTA models with much more parameters (e.g., Shikra-13B with 13B parameters) or much more training data (e.g., Qwen with 1.4B data). For VQA task, our model also shows significant advantages compared to other MLLMs. On VQAv2 val, dev and std, our model achieves 79.05, 81.04 and 81.17 accuracy respectively, which surpasses recent proposed Shikra with the same training data and procedure by a large margin, demonstrating the effectiveness of merging visual embeddings of DINOv2 and CLIP for enhancing visual capabilities. Besides, our COMM model outperforms Qwen with 1.54 and 0.58 accuracy gain on VQAv2 dev and OK-VQA respectively with less VQA training data, i.e., we use 0.6M and Qwen with 3.6M. Training with more VQA data might further improve performance and we leave it as future work.
5.5 Ablation Study
@@
Ablation on the MLP of DINOv2. We conduct ablation study on the design of the MLP module in DINOv2 for aligning visual and text embedding space. We ablate on the number and the expanding ratio of MLP module. Table 6 shows that increasing the number of MLP to 2 can evidently improve performance, demonstrate the effectiveness of using a more powerful network to align the vision only model DINOv2 to the word embedding space. However, increasing the number beyond 2 suffers the degraded performance. For the expanding ratio, increasing to 8 can improve performance, while increasing to 16 does not achieve significant performance gain. Moreover, we experiment with one linear layer, which suffers severe performance degradation. Thus, non-linear MLP is necessary for aligning the features of vision-only DINOv2 to the word embedding space.
Ablation on the visual model of MAE and DeiT. As shown in Table 7, MAE and DeiT suffers from evident performance degradation. For one thing, the visual features of MAE lack sufficient semantic information for global or regional understanding. For another, the supervised training of DeiT is so strong that it learns specialized visual space, making it difficult to align with the word embedding space.
@@
Table 7: Comparison of the visual model using CLIP, DINOv2 with our multi-level feature merging (MFM), MAE and DeiT. MAE-20 denotes using the features output by the 20-th layer of MAE. DeiT-20 denotes using the features output by 20-th layer.
Visual Model RefCOCO+ RefCOCOg RefCOCO POPE
test-A test-B val test-u val-u test-A test-B val A/P/R
CLIP w/ MFM 73.7 53.8 64.3 69.1 70.3 83.8 68.4 76.4 80.7/84.2/85.8
DINOv2 w/ MFM 75.3 59.3 67.0 73.0 71.8 84.4 74.1 79.6 80.3/84.2/85.5
MAE-20 64.7 49.4 56.8 63.7 62.8 77.9 68.6 73.6 66.8/71.1/76.7
MAE-22 65.9 50.0 58.5 64.2 63.2 79.3 69.8 74.9 68.0/71.2/77.5
DeiT-20 18.4 13.0 15.9 17.0 16.2 29.0 21.6 25.7 66.2/69.6/77.9
DeiT-22 25.3 15.4 19.4 22.6 21.8 36.9 25.3 32.0 67.9/71.6/78.7
5.6 Demonstrations
As shown in Fig. 7, our COMM model exhibits a multitude of promising capabilities including visual grounding, fine-grained region understanding and robustness to object hallucination. The first example showcases our strong fine-grained perception ability, which identifies implicit strawberries in a blender. The second example exhibits our strong visual grounding ability to successfully locates the object of sugar. The third case demonstrates our robustness to object hallucination. In contrast, Shikra fails on these challenging cases, showing the superior capabilities of our model. We provide additional demonstrations of our COMM model in this section to demonstrate a multitude of promising capabilities including visual grounding, fine-grained region understanding and robustness to object hallucination. For instance, we showcase Referring Expression Comprehension in Fig. 5 and Object Hallucination in Fig. 6.
@@ 5.6 Demonstrations
6 Conclusion